# Engagement Maximization

### v4

@@ 4 Implications
Implication 2: Commitment is Unnecessary.
Consider a modified version of our model in which the principal lacks commitment with respect to the law of the stochastic process $q_{t}$ and instead simply chooses signals for the agent to receive at each instant. In this case, we also assume that the agent lacks commitment with respect to her stopping rule. It is immediate that if both agents expect the policies of the solution with commitment to played going forward, they will be willing to adopt those policies at the current moment. [^17] That is, there are no profitable one-shot deviations from the commitment solution. To claim that as a result, the commitment solution is an equilibrium even in the absence of commitment, we would have to define sub-game perfection in the context of our continuous-time game.
The intuition behind this statement is that, under the optimal policy, the principal never promises the agent anything other than the minimum possible utility starting from any belief $q\in\mathcal{P}(X)$. That is, although it is in theory possible for a principal with commitment to induce the agent to continue by promising a larger-than-minimum level of utility in the future, such promises are not necessarily a feature of the optimal policy with commitment, and consequently commitment is unnecessary.
However, a lack of commitment can restrict the set of optimal policies. If beliefs change in a way that is strictly suspensive, $V_{R}(q_{t},\pi^{*})>0$, then the agent’s participation constraint becomes slack. The principal at this point would be tempted to re-optimize (inducing a different $\pi^{**}\neq\pi^{*})$. In our example, the set of suspensive beliefs was $[q_{0},1-q_{0}]$ for any $q_{0}<\frac{1}{2}$. But it is apparent from Figure 2 that if $q_{t}\in(q_{0},1-q_{0})$, the optimal posteriors given the initial belief $q_{t}$ are no longer in the support of $\pi^{*}$. Commitment is thus necessary for the agent to tolerate moving to strictly suspensive beliefs. In the absence of commitment, the only optimal policies are the kinds of dilutions described in Proposition 2 and policies that mix such dilutions with jumps to other equally suspensive beliefs (such as the policies illustrated in Figure 3).
Implication 3: Extreme Beliefs.
@@
See the appendix, section A.1. ∎
This proposition is illustrated in Figure 2. For all priors $\bar{q}_{0}\in(\underline{q},\bar{q})$, which is the continuation region in the benchmark model defined by $V^{B}(\bar{q}_{0})>\hat{u}(\bar{q}_{0})$, $q^{1}<\underline{q}<\bar{q}<q^{2}$, and consequently $Q^{p}(\bar{q}_{0})=\{q^{1},q^{2}\}\not\subset\hat{Q}^{a}(q_{0})=[\underline{q},\bar{q}]$.
Extreme beliefs are a natural consequence of the tradeoffs facing the principal. By forcing the agent to ultimately acquire more information than she would choose for herself (the extreme beliefs), the principal can simultaneously delay the agent’s stopping decision while providing information the agent is willing to attend to. Interpreted in the context of an internet platform and a user described earlier, our extreme beliefs result implies that the platform should provide in-depth content on a narrower set of topics than would be preferred by the user. We can interpret this loosely as encouraging the user to go down “rabbit holes” that lead to extreme beliefs.
Implication 4: The Necessity of Jumps in Beliefs.
@@
We have shown in Proposition 3 that, at all times under an optimal policy, beliefs are either suspensive or decisive. This result immediately implies that the optimal policy cannot involve continuous sample paths for beliefs (we assume the initial prior lies in the continuation region). If sample paths were continuous, the belief process $q_{t}$ would have to exit $\Delta(\pi^{*})$, which would immediately induce the agent to stop. Consequently, beliefs must jump discontinuously from the set of suspensive beliefs (which lie in the continuation region of the benchmark model) to the set of decisive beliefs, which lie in the strict stopping region of the benchmark model (by Proposition 4).

### v5

@@ 3.4 Implications of Engagement Maximization
Implication 2: Extreme Beliefs.
This result is described in Proposition 3 and is illustrated in Figure 2 below. Extreme beliefs are a natural consequence of engagement maximization. By forcing the agent to ultimately acquire more information than she would choose for herself (the extreme beliefs), the principal can simultaneously delay the agent’s stopping decision while providing information the agent is willing to attend to. Interpreted in the context of an internet platform and a user, our extreme beliefs result implies that the platform should provide in-depth content on a narrower set of topics than would be preferred by the user. We can interpret this loosely as encouraging the user to go down “rabbit holes” that lead to extreme beliefs. In the teacher-student context, extreme beliefs have a more benign interpretation: the teacher induces test-motivated students to learn more than they would choose to learn on their own.
3.5 Illustrative Examples
Example 1.
We illustrate our results thus far in a simple example. An agent faces a choice between two actions, $A=\{l,r\}$. The payoffs from the actions are uncertain and depend on the state of the world $x\in X=\{L,R\}$. The agent assigns equal prior probability to both states ($\bar{q}_{0}$ is uniform). The agent gets utility one when the chosen action matches the state ($u_{l,L}=u_{r,R}=1$) and utility negative one otherwise ($u_{l,R}=u_{r,L}=-1$). The agent is impatient and pays a constant cost of delay of two utils per unit of time ($\bar{\kappa}=2$). We assume that the information processing constraint $H$ is the negative of Shannon’s entropy, and will begin by assuming that the engagement measurement function $G$ is also proportional to Shannon’s entropy.
To build intuition, we will restrict attention to the following type of dilution signals (which is without loss, as this class includes the optimal signal structure). Each signal has two possible realizations $\{s_{l},s_{r}\}$, which arrives at Poisson rates that depend on the state (symmetrically):
L R
$s_{l}$ $\alpha\cdot p$ $\alpha\cdot(1-p)$
$s_{r}$ $\alpha\cdot(1-p)$ $\alpha\cdot p$
The information processing constraint for this class of signals requires that $\alpha$ and $p$ must satisfy
where $H^{S}(\cdot)$ is Shannon’s entropy. Note that $H^{S}\left(\frac{1}{2}\right)-H^{S}(p)$ is an increasing and convex function of $p$. Intuitively, both more frequent and more precise signals require processing “more information.” However, there is a trade-off between the frequency and the precision of signals.
We now explain why the agent is willing to wait for the arrival of the signal. If the agent has not observed either signal by time $t$, her beliefs will remain equal to her prior (and hence assign equal probability to each state). As a result, if the agent were to stop and act before observing either of the Poisson signals, her expected payoff would be zero. Because the problem is stationary, $V^{B}$ is also the agent’s expected utility at any time $t>0$, conditional on not having received either Poisson signal yet; it follows that the agent will strictly prefer waiting for the arrival of one of these signals to stopping. In fact, the results of Hébert and Woodford (2023) show that this signal structure and stopping strategy is optimal for the agent, both within the class of signals considered in this example and within the larger class of signals satisfying the mutual information bound. This strategy also yields a positive payoff for the principal. Because we have assumed that the measure of engagement is the negative of Shannon’s entropy, the expected engagement in this case is $\alpha^{-1}$ (which is also the expected decision time).
However, it is possible for the principal to do better. If the prinicpal exactly satisfies the agent’s information processing constraint (which will be optimal), within this class of Poisson signals, maximizing engagement requires increasing the expected decision time. This is only possible if the principal offers signals that are less frequent but more informative. Such signals are less desirable for the agent than the agent-optimal signals (by definition); but as established above, the agent strictly benefits from receiving the agent-optimal signals, and therefore might be willing receive somewhat less desirable signals as well.
Using these less frequent, more informative signals allows the principal to increase expected engagement ($\tilde{\alpha}^{-1}>\alpha^{-1}$) at the agent’s expense. It is straightforward in this example to see that using an even less frequent, even more informative signal would allow the principal to drive the agent to her indifference point while further increasing expected engagement. Note that utility is not transferable in this example (or in our general model); thus, the necessity of welfare minimization under the principal’s optimal strategy is not immediate from the setup of the model, and in fact will not hold when the principal is restricted to a subset of strategies (see Section 5).
We next illustrate Theorem 1 in the context of this example. We will now vary the agent’s prior belief $\bar{q}$, and focus our analysis on the agent’s posteriors when stopping ($q_{\tau}$) as opposed to the compound Poisson signals that generate those posteriors. Figure 1 illustrates the optimal $q_{\tau}$ for every prior belief: When $\bar{q}_{0}\leq\underline{q}$ or $\bar{q}_{0}\geq\bar{q}$, the agent stops immediately no matter what information she will receive. When $\bar{q}_{0}\in(\underline{q},\bar{q})$, the dashed lines illustrate the support of the agent-optimal policy. The lines are flat, meaning that the agent’s stopping beliefs do not change with the prior belief, consistent with the prediction of the standard RI models. [^16] This property is known as the “locally invariant posteriors” (Caplin et al. (2022)). The solid curves illustrate the support of the unique optimal $\pi^{*}$. They are closer to the boundaries zero and one, indicating that the posterior beliefs become more polarized under the principal’s optimal policy.
Figure 1: $\text{Supp}(q_{\tau})$ as a correspondence of $q_{0}$
Figure 2: Engagement as a function of $q_{0}$
Example 2: Leading Cases.
We next adapt example 1 to illustrate our two leading cases, in the context of engaging test-motivated students. Suppose the true state of the world $T$ consists of both a test-relevant and test-irrelevant dimension, $T=T_{1}\times T_{2}=\{L,R\}\times\{0,1\}$. The student and teacher know that the student will be asked a single question, $Q=\{Q_{0}\}$, whose responses are $A=\{l,r\}$, with $Q_{0}(L0\}=Q_{0}(L1)=l$ and $Q_{0}(R0)=Q_{0}(R1)=r$. That is, the $\{L,R\}$ component of the true state is relevant, and the $\{0,1\}$ component is irrelevant. The student (agent) again faces a binary choice $A=\{l,r\}.$ The student’s utility is one if she answers the question correctly and negative one otherwise, and therefore is identical (ignoring the test-irrelevant dimension) to that in example 1. The prior belief is uniform. We continue to assume the agent’s information processing capacity is defined by the negative of Shannon’s entropy, and continue to use the parameters $\bar{\kappa}=2,\chi=1$.
Leading Case: $H=G$.
We first study the leading case of $H=G$, and illustrate Proposition 4 with a specific example. By Proposition 1, the teacher’s optimal strategy maximizes
$\mathbb{E}^{\pi}[\hat{u}(q)-(2-\lambda)(H(q)-H(\bar{q}_{0}))]$
for some $\lambda<2.$ The student-optimal benchmark maximizes
$\mathbb{E}^{\pi}[\hat{u}(q)-2(H(q)-H(\bar{q}_{0}))].$
Both optimization problem can be written as:
$-\sum_{a,t_{1}}\frac{1}{2}P(a|t_{1})\log(\sum_{t_{1}}\frac{1}{2}P(a|t_{1}))),$
which is equivalent to a static rational inattention problem. [^17] This equivalence is special to mutual information and is the implication of a more general “compression invariance” property introduced by Caplin et al. (2022); Bloedel and Zhong (2020). As shown in Proposition 4, when there is a test-irrelevant state that is costly to learn about, no information about the state will be acquired either in the student-optimal benchmark or in the principal-agent problem, even though the student learning about state $t_{2}$ enters the teacher’s payoff function.
R0R1L0L1Prior beliefAgent-optimal posteriorsPrincipal-optimal posteriors
Figure 3: Leading case: $H=G$.
Figure 3 illustrates the analysis above. The tetrahedron depicts the probability simplex that lies in $\mathbb{R}^{3}$. Each vertex of the tetrahedron is a degenerate belief (labeled by the state that occurs with probability one). The red dot is the uniform prior. The hyperplane represents the cutoff beliefs where the student is indifferent between actions $l$ and $r$. To the right of the hyperplane, $q_{R}>q_{L}$ and $r$ is optimal. Vice versa for the left of the hyperplane. The two blue dots are the student-optimal posterior beliefs. The two green dots are the solution to the principal-agent problem. Evidently, all the dots are on a line segment, which illustrates that the teacher provides only test-related information to the student. It is also evident that the teacher provides more information than the student would choose for herself.
Leading Case: Decision-Irrelevant $G$.
Now, we turn to the second leading case, in which the teacher only cares about how much the student knows about the state $t_{2}\in T_{2}$ (the test-irrelevant dimension). We assume that the teacher’s payoff is $2\mathbb{E^{\pi}}[|q_{t_{2}}-0.5|]$ if the measure of stopping belief is $\pi$, where $q_{t_{2}}$ denotes the probability of $t_{2}=1$ under the student’s stopping belief. [^18] This objective does not satisfy our differentiability assumptions for $G$ at all points, but the issues arising from non-differentiability will not be relevant for this example. Note that this payoff function is equivalent to the instrumental value of information in a hypothetical binary decision problem where the utility of matching (mismatching) the state $y$ is 1 (-$1$). Proposition 1 can be extended to show that the teacher’s optimal policy maximizes the following auxiliary problem (for some $\lambda>0$):
Per Matêjka et al. (2015), the solution is logit, given by
is identical to the student-optimal solution (and independent to $\lambda$). This is an illustration of the conclusion of Proposition 5. The unknown parameter $\lambda$ can be pinned down by setting the student’s IC condition binding. The analysis above suggests that when the teacher only cares about the test-irrelevant knowledge, she will give the student exactly his preferred test-relevant information and extra knowledge such that the student gets barely enough welfare to be willing to participate.
R0R1L0L1Prior beliefAgent-optimal posteriorsPrincipal-optimal posteriors
Figure 4: Leading Case: Decision-Irrelevant $G$.
4 Characterization of dynamics
In Section 3, we showed that to solve for the optimal policy, it is sufficient to study a simple static rational inattention problem (R) and implement the dynamic policy by diluting the static solution. In this section, we seek to find necessary conditions that characterize the dynamics of the optimal policy. For the purpose of characterizing optimal polices, for this section, we focus on the non-degenerate case where the optimal policy is to acquire some information ($\bar{J}(\bar{q}_{0})>0$ in the context of Proposition 1). We show that two desiderata—interim incentive compatibility and credibility—jointly provide a tight characterization of the belief dynamics.
4.1 Implications of interim incentive compatibility
To begin, we define, given a measure $\pi\in\mathcal{P}(\mathcal{P}(X))$, the maximal value of information acquisition in a hypothetical restricted static rational inattention problem:
$\overline{V}_{R}(q,\pi)=\max_{\pi^{\prime}\in\mathcal{P}(\text{Supp}(\pi)):E^{\pi^{\prime}}[q^{\prime}]=q}\mathbb{E}^{\pi^{\prime}}\left[\widehat{\mu}(q^{\prime})-\hat{u}(q)-\frac{\bar{\kappa}}{\chi}(H(q^{\prime})-H(q))\right].$
This problem maximizes the agent’s expected utility subject to the usual constraint of Bayes consistency and the additional constraint that the posteriors lie in the support of $\pi$. We use the notation $\overline{V}_{R}(q,\pi)$ to emphasize that the problem is restricted relative to the usual rational inattention problem. Note that the problem is feasible only for $q$ that lie in in the convex hull of the support of $\pi$, which we denote by $\text{Conv}(\text{Supp}(\pi))$.
We interpret $\overline{V}_{R}(q,\pi)$ as describing the maximum possible value of information acquisition, in the sense that $\overline{V}_{R}(q,\pi)$ is the difference between the best possible utility the agent could achieve with and without information acquisition, under the restriction that the set of stopping beliefs lie in the support of $\pi$ [^19] The value of information acquisition is related to, but not the same as, the notion of uncertainty defined in Frankel and Kamenica (2019). We define the set of “suspensive” beliefs given $\pi$ as the set for which the value of information acquisition is weakly positive.
Definition 2.
Given $\pi\in\mathcal{P}(\mathcal{P}(X))$, the belief $q\in\mathcal{P}(X)$ is suspensive if $q\text{$\in$Conv}(\text{Supp}(\pi))\setminus\mathrm{Supp(\pi)}$ and satisfies $\overline{V}_{R}(q,\pi)\geq 0$; the belief $q\in\mathcal{P}(X)$ is decisive if $q\in\text{Supp}(\pi)$.
Given a measure $\pi$ that describes the stopping beliefs, it is intuitive that beliefs in $\mathrm{Supp(\pi)}$ are “decisive” because they directly lead to stopping. At the suspensive beliefs, the agent could benefit, relative to the prior consistent with $\pi$, from the signals that the principal is offering (which always result in some posterior in the support of $\pi$). Therefore, suspensive beliefs are those that the principal can possibly induce “suspense” before eventually leading the agent to stopping beliefs in $\mathrm{Supp(\pi)}$.
Suppose the optimal policy in (R), $\pi^{*}$, is unique. In this case, the stochastic process for beliefs $q_{t}$ cannot leave the set of suspensive beliefs before $t=\tau$. If $q_{t}$ left the convex hull of the support of $\pi^{*}$, then $\pi^{*}$ would not describe the law of $q_{\tau}$. If $q_{t}$ was decisive or if $\overline{V}_{R}(q_{t},\pi^{*})<0$, the agent would choose to stop. In the case of multiple optimal policies, this argument applies for some optimal $\pi^{*}$, which leads to the proposition below. Note that this proposition and the one that follows will include a technical caveat about finite support (that we believe is inessential), to simplify the proofs that follow. [^20] The assumption of finite support avoids the need to carefully define conditional expectations.
@@ 4.1 Implications of interim incentive compatibility
Given $\bar{q}_{0}\in\mathcal{P}(X)$, let $\Pi^{*}$ be the set of solutions to (R). Suppose $(P,I^{H},\tau^{*})$ solves the principal’s problem under full commitment (Definition 1) and $\text{Supp}(q_{t}|t<\tau)$ has finite support for all $t$, then, $\forall t\in[0,\tau)$, $\forall q\in\text{Supp}(q_{t}|t<\tau)$, $q$ is suspensive given some $\pi$:
$\sup_{\pi\in\Pi^{*}}\overline{V}_{R}(q,\pi)\geq 0.$
@@
By Theorem 1, there is some $\pi^{*}\in\Pi^{*}$ that is the law of $q_{\tau^{*}}$ under the optimal policies, and the second claim follows. By the martingale property of the belief process, $q_{t}=E^{P^{*}}[q_{\tau}|\mathcal{F}_{t},t<\tau^{*}]$ for all $t\in[0,\tau^{*})$, and thus $q_{t}$ must lie in the convex hull of the support of $\pi^{*}$. The fact that $t<\tau^{*}$ implies $J(t)>0$ (i.e. $q_{t}\not\in\text{Supp}(\pi^{*})$), as otherwise, the principal cannot induce the agent to continue. Let $\pi^{\prime}$ be the distribution of $q_{\tau^{*}}$ conditional on $q_{t},t<\tau^{*}$. The agent’s incentive compatibility implies $\frac{\bar{\kappa}}{\chi}\mathbb{E}^{\pi^{\prime}}[H(q)-H(q_{t})]\leq\mathbb{E}^{\pi^{\prime}}[\widehat{\mu}(q)-\widehat{\mu}(q_{t})]$, from which it follows immediately that $\overline{V}_{R}(q_{t},\pi^{*})\geq 0$. ∎
Solving (R) generically leads to the IC constraint binding, which suggests that the value of acquiring $\pi^{*}$ is exactly zero at the prior for the agent. The set of suspensive beliefs thus defines beliefs that are “more uncertain” than the prior. Therefore, the interpretation of Proposition 6 is that two types of signals appear in any optimal policy almost surely. The first type are decisive signals that are so informative that the agent makes a decision immediately after receiving the signal. The second type are suspensive signals that causes the agent’s beliefs to move to a posterior that is “more uncertain” than the prior.
Implication 3: The Necessity of Jumps in Beliefs.
@@
We have shown in Proposition 6 that, at all times under an optimal policy, beliefs are either suspensive or decisive. This result immediately implies that the optimal policy cannot involve continuous sample paths for beliefs (we assume the initial prior lies in the continuation region). If sample paths were continuous, the belief process $q_{t}$ would have to exit the set of suspensive beliefs, which would immediately induce the agent to stop. Consequently, beliefs must jump discontinuously from the set of suspensive beliefs (which lie in the continuation region of the benchmark model) to the set of decisive beliefs, which lie in the strict stopping region of the benchmark model (by Proposition 3).
4.2 Implications of credibility
So far, our analysis assumes full intertemporal commitment of the principal. However, in practice, commitment power might be limited, leading to a credibility concern. In this subsection, we tackle the issue of limited commitment by characterizing the necessary and sufficient conditions for a policy to be credible, i.e., immune to deviations in a sub-game perfection sense.
The definition of sub-game perfection for a continuous time game like ours is technically delicate. To avoid these complexities, we introduce the notion of a "strongly sub-game perfect" equilibrium. An equilibrium is strongly sub-game perfect if the principal is not willing to deviate, even if he can deviate to the full-commitment solution, and likewise the agent is willing to stop even if she can instead switch to the agent-optimal benchmark at that point.
In an environment in which commitment is valuable, such an equilibrium cannot exist. We will show in our environment that some (but not all) full-commitment solutions to the principal-agent problem are strongly sub-game perfect in this sense, which is to say that commitment has no value. Nevertheless, despite this lack of value, the principal’s inability to commit will prevent some full-commitment solutions from being implemented as equilibria of the game without commitment.
Definition 3.
$\forall q_{0}\in\mathcal{P}(X)$, $(P,I^{H},\tau)\in\mathcal{A}(q_{0})$ constitutes a strongly sub-game perfect equilibrium of the game if the agent’s interim incentive compatibility constraints are satisfied and
  1. (i)
    the principal would not deviate to the commitment solution: $\forall t$ and $q\in\text{Supp}(q_{t}|t\leq\tau)$, $\mathbb{E}^{P}[G(q_{\tau})-G(q_{t})|\mathcal{F}_{t},q_{t}=q]\geq J(q)$, and
  2. (ii)
    the agent would not deviate by continuing under the agent-optimal benchmark: $\forall q\in\text{Supp}(q_{\tau})$, $V^{B}(q)\leq\hat{u}(q)$.
Let $\mathcal{V}:\mathcal{P}(X)\rightrightarrows\mathbb{R}_{+}^{2}$ denote the correspondence from prior to the set of strongly sub-game perfect equilibrium payoffs of the principal and agent.
This notion of strongly sub-game perfect equilibrium is “stronger” than sub-game perfection in the sense that we overestimate deviation payoffs in conditions (i) and (ii): upon a deviation at belief, we award the principal with the full commitment payoff and the agent with her agent-optimal payoff, respectively. If such an equilibrium exists, it is clearly sub-game perfect in the usual sense as well.
In the analysis that follows, we will call a full-commitment optimal policy $(P,I^{H},\tau)$ credible if $(P,I^{H},\tau)$ is also a strongly sub-game perfect equilibrium in the game without commitment. The following theorem shows that the optimal dilution policy we construct in section 3 is indeed credible and hence constitutes a (strongly) sub-game perfect equilibrium. To avoid technicalities, what we will not discuss is whether there are other sub-game perfect equilibria that involve lower payoffs for the principal.
Theorem 2.
$\forall\bar{q}_{0}\in\mathcal{P}(X)$, $(J(\bar{q}_{0},0))\in\mathcal{V}(\bar{q}_{0})$. Suppose $(P,I^{H},\tau)$ solves the principal’s problem under full commitment (Definition 1) and $\text{Supp}(q_{t}|t<\tau)$ has finite support for all $t$. $(P,I^{H},\tau)$ is credible if (and satisfies condition (i) of a strongly sub-game perfect equilibrium only if) the agent’s interim surplus is 00:
$\forall t\text{ and }q\in\text{Supp}(q_{t}|t<\tau),\quad\mathbb{E}^{P}[\hat{u}(q_{\tau})-\hat{u}(q)-\bar{\kappa}(\tau-t)|t<\tau]=0.$ (4)
Proof.
See Appendix section B.1. ∎
Note since $(J(\bar{q}_{0}),0)\in\mathcal{V}(\bar{q}_{0})$, $J(\bar{q}_{0})$ is itself a strongly sub-game perfect equilibrium payoff. Condition (i) in 3 must be satisfied for any weaker equilibrium notion that “selects SPNE in all sub-games in favor of the principal.” Therefore, (4) is a sufficient and necessary condition that guards the principal from deviations followed by a favorable selection of continuing equilibrium.
Using this condition, we can characterize a set of beliefs that cannot be reached in any credible solution to the principal-agent problem. To do this, we will define the set of “strictly suspensive” beliefs. In parallel to our definition of suspensive beliefs, we start by defining, given a measure $\pi\in\mathcal{P}(\mathcal{P}(X))$, the minimal value of information acquisition in a hypothetical restricted static rational inattention problem:
$\underline{V}_{R}(q,\pi)=\min_{\pi^{\prime}\in\mathcal{P}(\text{Supp}(\pi)):E^{\pi^{\prime}}[q^{\prime}]=q}\mathbb{E}^{\pi^{\prime}}\left[\widehat{\mu}(q^{\prime})-\hat{u}(q)-\frac{\bar{\kappa}}{\chi}(H(q^{\prime})-H(q))\right].$
We will say that a belief is strictly suspensive if this value is strictly positive.
Definition 4.
Given $\pi\in\mathcal{P}(\mathcal{P}(X))$, the belief $q\in\mathcal{P}(X)$ is strictly suspensive if $q\ \text{ $\in$ Conv}(\text{Supp}(\pi))\setminus\mathrm{Supp(\pi)}$ and satisfies $\underline{V}_{R}(q,\pi)>0$.
Note the contrast between suspensive beliefs and strictly suspensive beliefs. A belief is suspensive given $\pi$ if the maximum value of information is weakly positive, but is strictly suspensive only if the minimum value of information is strictly positive.
When $\pi^{\prime}=\pi$ is the only measure on the support of $\pi$ that satisfies Bayes-consistency, there is no distinction between $\bar{V}_{R}$ and $\underline{V}_{R}$. [^21] This will be true generically when the number of actions is weakly lower than the number of states. In this case, a belief is suspensive (resp. strictly suspensive) when this function is weakly (resp. strictly) positive.
The following proposition demonstrates that a solution to the principal-agent problem is credible only if interim beliefs are not strictly suspensive.
Proposition 7.
Given $\bar{q}_{0}\in\mathcal{P}(X)$, let $\Pi^{*}$ be the set of solutions to (R). Suppose $(P,I^{H},\tau)$ solves the principal’s problem under full commitment (Definition 1) and $\text{Supp}(q_{t}|t<\tau)$ has finite support for all $t$. $(P,I^{H},\tau)$ is credible only if $\forall t\in[0,\tau)$, $\forall q\in\text{Supp}(q_{t}|t<\tau)$, $q$ is not strictly suspensive given some $\pi$:
$\inf_{\pi\in\Pi^{*}}\underline{V}_{R}(q,\pi)\leq 0.$
Proof.
By Theorem 1, there is some $\pi^{*}\in\Pi^{*}$ that is the law of $q_{\tau^{*}}$ under the optimal policies. By the martingale property of the belief process, $q_{t}=E^{P^{*}}[q_{\tau}|\mathcal{F}_{t},t<\tau^{*}]$ for all $t\in[0,\tau^{*})$, and thus $q_{t}$ must lie in the convex hull of the support of $\pi^{*}$. Let $\pi^{\prime}$ be the distribution of $q_{\tau^{*}}$ conditional on $(q_{t},t<\tau^{*})$. Theorem 2 implies $\frac{\bar{\kappa}}{\chi}\mathbb{E}^{\pi^{\prime}}[H(q)-H(q_{t})]\geq\mathbb{E}^{\pi^{\prime}}[\widehat{\mu}(q)-\widehat{\mu}(q_{t})]$, from which it follows immediately that $\underline{V}_{R}(q_{t},\pi^{*})\leq 0$. ∎
Combining Propositions 6 and 7, in any solution that satisfies interim incentive compatibility and credibility, interim beliefs are suspensive but not strictly suspensive. They must be suspensive, as otherwise the agent would choose to stop, but cannot be strictly suspensive, as otherwise the principal would be tempted to deviate and extract more of the surplus.
Implication 4: Credible Beliefs Cannot Diffuse.